OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, -2). (4 points) Choices: y = 2x + 4 y = -1/2x - 1/2 y = − 1/2 x - 7/2 y = 2x - 8 My answer was: y = 2x + 4 I just want to know if I was correct. :)

OpenStudy (steve816):

First of all, remember that parallel lines have the same slope, so the choices B and C are knocked out.

OpenStudy (steve816):

Then, you want to plug in those coordinates (3,-2) to the point-slope form. \[y-(-2)=2(x-3) \] \[y+2 = 2x-6\] \[y=2x-8\] So D would be the correct answer :)

OpenStudy (steve816):

Hope this helps my friend!

OpenStudy (anonymous):

@steve816 This helps a lot! Thank you so much. :D Happy studying. :) (Sorry, I'm new here, but how do I award medals? )

OpenStudy (steve816):

I'm glad I was able to help! To award medal, simply press on the best response button next to the medal. Cheers!

OpenStudy (anonymous):

Ah, gotcha! Thank you again. Cheers! :)